OpenStudy (anonymous):

HEEEELP!! :( there were 22 guest. the host chooses 14 people to sit around a circular table and the others around another circular table. what is the probability that simon and brenna are not seated on the same table?

OpenStudy (anonymous):

i've been sooo confused! i understand that for one table it's c(22,14) so that leaves the other one with c(8,8) possibilities? did i get that part right?? then i know there's a technique wherein total number of ways-total number of ways together=total number of ways apart but i can't seem to apply it to this! :(

OpenStudy (anonymous):

@ParthKohli ? @sauravshakya ?

OpenStudy (anonymous):

see, the total no. of ways in which 22 people can seat on two round tables in set of 14 and 8 are 13! * 7! Now, the catch here is to find out the no. of ways in which simon and brenna can seat together: If simon and brenna sit on table with 14 people or they can on the table with 8 people. The total no. of ways in which they can sit on same table is, 13! +7! So, the probability = (13! +7!) / (13! * 7!)..

OpenStudy (anonymous):

@bhaweshwebmaster don't we need combination in order to choose 14 out of the 22 and all that?

OpenStudy (anonymous):

@Hero ? @TuringTest ? @saifoo.khan ?@Mertsj ? anyone?

OpenStudy (anonymous):

@dydlf No Sir, we needn't use combination for this. This is the problem from "Basic Principles of Counting"... No need of combination...

OpenStudy (anonymous):

but there are different ways of getting those 14 people out of the 22 :(

OpenStudy (anonymous):

I think you are looking at wrong direction. factorial must do it...

OpenStudy (anonymous):

@bhaweshwebmaster wait... why just factorial, then? because i could get a b c d e f g h i j k l m n o p q r s t u v -- 22 people and get a b c d e f g h i j k l m n as my 14 people or i j k l m n o p q r s t u v or even e f g h i j k l m n o p q r that's why i think combination is needed... @amistre64 ? @Hero ? @ParthKohli ? anyone? what do you guys think?

OpenStudy (anonymous):

Ya I understood what you were trying to say.. and probably you are right.. (or probably not).. I am 50-50... perhaps, math genuises on the forum could give us an insight better..

OpenStudy (anonymous):

thanks @bhaweshwebmaster ! i'm really lost here. hopefully someone could help me! :(

OpenStudy (amistre64):

im wondering how much of the information given is just for distraction

OpenStudy (amistre64):

what is the probability that they are seated at the same table might be simpler to construct, then we can take the complement is one idea i have

OpenStudy (amistre64):

there are 4 outcomes that i see; using table A and B and persons s and b, the possible seating outcomes are: As, Ab As, Bb Bs, Ab Bs, Bb The possibility of you walking into the room and seeing s and b sitting at the same table would then be 2/4, and the probability that they are not seated at the same table is 2/4.

OpenStudy (amistre64):

the possbility that s is at table A is say ... 1/14 the possbility that s is at table B is say ... 1/8 and same notions for person b if you wanna try to construct a tree chart, but im not sure what all you have been studying to know what method they are trying to get you to work out

OpenStudy (amistre64):

another idea is: what is the probability of s being picked 1st, 2nd, .... 14th ? and b being picked 15th, 16th, ... 22nd?

OpenStudy (amistre64):

P(s=1) = 1/22 P(s=2) = 1/21 P(s=3) = 1/20 ... \[P(s=1\to14)=\sum_{k=9}^{22}\frac{1}{k}\] just wondering :)

OpenStudy (anonymous):

Isnt it {2! 20C13} / {22C14}

OpenStudy (anonymous):

@sauravshakya why 2! 20C13 for the numerator? can you explain please :D

OpenStudy (anonymous):

Because simon and brenna are not seated on the same table

OpenStudy (anonymous):

So, at first they must sit on different tables [ which mean 2!] And we have to choose 13 people out of remaining 20 people to sit in the table [which gives 20C13]

OpenStudy (anonymous):

So, 2! * 20C13 for the numerator

OpenStudy (anonymous):

makes sense! thank you!!

OpenStudy (anonymous):

welcome